the father was 8 times older than his doughter 4 years ago,in twenty years time the father wil be twice the age of the doughter. 1)how old is the father now? 2)how old wl be the doughter in 30 years time?

Answer by Maths68(1474)

Let
Present age of Father = f
Present age of Daughter = d
4 years ago
Age of father = f-4
Age of daughter = d-4
Then
The father was 8 times older than his daughter
f-4=8(d-4)
f-4=8d-32
f=8d-32+4
f=8d-28......................(1)

In twenty years time
Age of father will be = f + 20
Age of daughter will be = d+ 20
Then
father will be twice the age of the daughter.
f+20=2(d+20)
f+20=2d+40
f=2d+40-20
f=2d+20.......................(2)

Plug in the value of f from (1) to (2)
8d-28=2d+20
8d-2d=20+28
6d=48
6d/6=48/6
d=8
Put the value of d in (1)
f=8d-28
f=8(8)-28
f=64-28
f=36


Present age of Father = f = 36 years old
Present age of Daughter = d = 8 years old
Daughter will be 38 years old after 30 years.